OpenStudy (johnny1423):

An exercise gives the parameterized equation of a movement ( r(t)=x(t),y(t),z(t) ) ,and asks to prove that the movement is happening on one plane and that it is circular. How do I do that?

OpenStudy (michele_laino):

what are the functions: \(x(t),y(t),z(t)\) ?

OpenStudy (johnny1423):

x=2+(cost/\sqrt(2))+(sint/\sqrt3), y=2-(cost/\sqrt2)+(sint/\sqrt3), z=1+(sint/\sqrt3)\]

OpenStudy (michele_laino):

a simple computation, shows that the subsequent condition holds: \[x + y - 2z = 2\] please check, by substitution. Such equation, in the euclidean space, is the cartesian equation of a plane

OpenStudy (michele_laino):

if we differentiate the coordinates: \(x,y,z\), we get: \[\left\{ \begin{gathered} \dot x = \frac{{ - \sin t}}{{\sqrt 2 }} + \frac{{\cos t}}{{\sqrt 3 }} \hfill \\ \hfill \\ \dot y = \frac{{\sin t}}{{\sqrt 2 }} + \frac{{\cos t}}{{\sqrt 3 }} \hfill \\ \hfill \\ \dot z = \frac{{\cos t}}{{\sqrt 3 }} \hfill \\ \end{gathered} \right.\]

OpenStudy (michele_laino):

so the tangent vector \(T\), is: \[T = \left( {\frac{{ - \sin t}}{{\sqrt 2 }} + \frac{{\cos t}}{{\sqrt 3 }},\;\;\frac{{\sin t}}{{\sqrt 2 }} + \frac{{\cos t}}{{\sqrt 3 }},\;\;\frac{{\cos t}}{{\sqrt 3 }}} \right)\] furthermore, we have: \[{\left( {ds} \right)^2} = {\left( {dx} \right)^2} + {\left( {dy} \right)^2} + {\left( {dz} \right)^2} = {\left( {dt} \right)^2},\quad \Rightarrow ds = dt\]

OpenStudy (michele_laino):

next, I apply the formulas of \(Differential\;\;Geometry\), and I get: \[\frac{{dT}}{{ds}} = \frac{{dT}}{{dt}} = \dot T = \left( {\frac{{ - \cos t}}{{\sqrt 2 }} - \frac{{\sin t}}{{\sqrt 3 }},\;\;\frac{{\cos t}}{{\sqrt 2 }} - \frac{{\sin t}}{{\sqrt 3 }},\;\;\frac{{ - \sin t}}{{\sqrt 3 }}} \right)\]

OpenStudy (michele_laino):

and I can write this: \[\left\| {\dot T} \right\| = 1,\quad \Rightarrow r = \frac{1}{{\left\| {\dot T} \right\|}} = 1\] namely the radius of curvature of trajectory, is \(constant\)

OpenStudy (johnny1423):

Thanks!